OpenStudy (anonymous):

Anyone? Let f(x) = x + 2 and g(x) = x2 – 6x + 3. Find g(f(x)). x2 – 6x + 5 x2 – 5x + 5 x2 – 2x – 5 x2 + 10x + 19

OpenStudy (jdoe0001):

think of it this way g( f(x) ) ^ g is HOSTING f(x), who is inside

OpenStudy (jdoe0001):

so any "x" in the HOST function, will be replace by the GUEST function

OpenStudy (anonymous):

wait what? I'm a little bit confused.

OpenStudy (jdoe0001):

$$ f(x) = \color{red}{x+2}\\ g(x) = x^2-6x+3\\ g ( f(x) ) = (\color{red}{x+2})^2-6(\color{red}{x+2})+3 $$

OpenStudy (jdoe0001):

so, expand that, cancel like-terms, and you'd get g(f(x))

OpenStudy (anonymous):

Well, f(x) = x + 2, so we plug that in for the x values of g(x), so: g(f(x)) = (x + 2)^2 - 6(x + 2) + 3 = x^2 + 4x + 4 - 6x - 12 + 3 = x^2 - 2x - 5

OpenStudy (anonymous):

Okay thank you. I get it now.

OpenStudy (jdoe0001):

yw